OpenStudy (anonymous):

how to graph general rational functions? like y=2x/x^2-1

OpenStudy (anonymous):

first, you'll need to know the asymptotes. in this case, as\(x^2 -1 \neq 0\) \(x= \pm 1\) is the asymptotes.

OpenStudy (anonymous):

then you'll need to know the max/min points. use differentiation to find them and plot the points.

OpenStudy (anonymous):

so like two and 1? also how do you know or like get if it has a 0 or not?

OpenStudy (anonymous):

\(y=2x(x^2-1)^{-1}\) \(y'=2(x^2-1)^{-1} -2x(x^2-1)^{-2} \)

OpenStudy (anonymous):

how did you get the two and 1?

OpenStudy (anonymous):

umm cuz there's a 2 and a 1? i was just trying but i knew it probably wasn't that easy

OpenStudy (anonymous):

nope. you'll have to use the global max/min test to find the points. at the max points, their gradients=0, so y'=0.

OpenStudy (anonymous):

is it because it's 2x it's zero? like cuz of the x or if not then plz explain that that global max/min test is cuz i don't get it

OpenStudy (anonymous):

at the max point, 

OpenStudy (anonymous):

wait. is this a precalculus question?

OpenStudy (anonymous):

? that looks higher than 0 aka the 0 on the y axis unless that has somethin to do with the gradient cuz idk what that is and yes this is a precalc question though my algebra book has it in there.

OpenStudy (anonymous):

I put the question here and like in a few other math places cuz i thought maybe someone would answer my question cuz when i put it in only one i was waiting for like ten minutes and nobody answered

OpenStudy (anonymous):

okay, then you might want to know that it should look like this. without calculus, you'll just have to memorise the shape and plot points.